OpenStudy (anonymous):

when solving a radical equation john and tim came to 2 different conclusions john found a solution while tim solutions did not fit the equation Create an justify 2 situations were john is correct and antheir where tim is correct.

OpenStudy (anonymous):

@thomaster

OpenStudy (amoodarya):

\[\sqrt{x}-x=0\] \[\sqrt{x}=x \rightarrow (\sqrt{x}=x)^2\] \[x=x^2 \rightarrow x-x^2=0 \rightarrow x(1-x)=0 \rightarrow x=0 , x=1\] both of roots are correct if check

OpenStudy (amoodarya):

another equation with the same idea \[\sqrt{x}+x=0\] \[\sqrt{x}=-x \rightarrow (\sqrt{x}=-x)^2\] \[x=x^2 , x(1-x)=0 \rightarrow x=0 , x=1\] x=0 is correct but x=1 is not because sqrt(1)+1=1+1=2 instead of 0